OpenStudy (surryyy):

The point (3, 0) is a solution of the equation: y = 2x − 11 Answer True or False

OpenStudy (anonymous):

This you can check yourself, can't you?

OpenStudy (anonymous):

You are given \(x\) and you are given \(y\) too, just plug in the equation given, and check whether right hand side equals left hand side or not?

OpenStudy (anonymous):

Am I talking to someone here or with myself only?? :)

OpenStudy (surryyy):

it would be false right?

OpenStudy (anonymous):

Do you have keyboard? @surryyy

OpenStudy (anonymous):

If it is false, then how can it be Right?? :P

OpenStudy (anonymous):

yep, you are correct.. It does not satisfy the equation.. :)

OpenStudy (anonymous):

Good..

OpenStudy (surryyy):

k Thnx lol

OpenStudy (anonymous):

\((3,0)\), whenever you are given point like this, the first value ie \(3\) indicates x-coordinate and second value ie \(0\) indicates y-coordinate.. So, you have \(x\) and \(y\) values with you, \(x=3\) and \(y=0\)..

OpenStudy (anonymous):

Now, your equation is : \(y = 2x - 11\) Plug the values in: \(0 = 2(3) -11\) \(0 = 6 - 11\) but \(0 \ne -5\), so LHS is not equal to RHS..

OpenStudy (anonymous):

As this solution is dissatisfying the equation, so it is not its solution.. :)

OpenStudy (anonymous):

\(\text{This is it..}\)

OpenStudy (anonymous):

Wow...really informative. Thanks @waterineyes you make me understand more about this lesson.

OpenStudy (anonymous):

you are welcome dear.. :)

OpenStudy (anonymous):

Thanks for ur reply...It made my day :)